The mean gasoline consumption of 10 cars is 28 liters with a standard

deviation of 1.6 liters. Find the interval estimate using 95% confidence

interval.



1
Expert's answer
2022-04-04T15:58:25-0400

"95\%CI=(\bar x-t_{0.025,n-1}\frac{s}{\sqrt{n}},\bar x+t_{0.025,n-1}\frac{s}{\sqrt{n}})="

"=(28-2.26\frac{1.6}{\sqrt{10}},28+2.26\frac{1.6}{\sqrt{10}})=(26.86,29.14)."